write an equation for the line with slope 1/3 and y - intercept 4

Found 2 solutions by nerdybill, richwmiller:
Answer by nerdybill(7384)

write an equation for the line with slope 1/3 and y - intercept 4
.
"Slope-intercept" form:
y = mx + b
where
m is the slope
b is the y-intercept
.
Plugging in the given data, we have:
y = (1/3)x + (-4)
y = (1/3)x - 4

Answer by richwmiller(17219)